A gas takes up a volume of 17 L, has a pressure of 2.3 atm, and a temperature of 299 K. If the temperature raised to 350 K and pressure is lowered at 1.5 atm, what is the new volume of gas?




1
Expert's answer
2021-06-07T03:12:45-0400

"P_1=2.3 atm"

"V_1=17 l"

"T_1=299 K"


"P_2=1.5 atm"

"T_2=350 K"

Let "V_2=V_2"


Now:

"\dfrac{P_1V_1}{T_1} = \dfrac{P_2V_2}{T_2}"


"V_2= \dfrac{2.3\times17 \times350}{1.5\times299}=30.51l"